Falconmaster:

If D varies directly with t and D is 105 when t is 3, find D when t is 8. 35 280 840

Falconmaster:

@Ultrilliam

celticcat:

Direct variation means that D = kt where k is a constant to find k plug in the given values of D and t:- 105 = k*3 k = 105 / 3 = 35 o we have D = 35t when t = 8 D = 35*8 = 280